OpenStudy (theprofessor35):

VOLUME QUESTION Jayden is making oatmeal cookies for a bake sale. He needs 3 cups of oatmeal for each batch of cookies. He has a container of oatmeal shaped like a cylinder, with a radius of 6 cm and a height of 24.5 cm. The 1 cup measuring cup is also shaped like a cylinder and has a radius of 3.9 cm and a height of 5 cm. How many batches of oatmeal cookies can Jayden make?

OpenStudy (ace-n-it):

He needs 3 cups of oatmeal for each batch of cookies.

OpenStudy (theprofessor35):

Yes....

OpenStudy (theprofessor35):

I'm not sure. Because this is a volume question, I would do the formula thing and idek. It's fine, I'll wait for someone to help :)

OpenStudy (xmissalycatx):

I can help you. :P Do you know the volume formula for a cylinder?

OpenStudy (theprofessor35):

Uh...I think it's v = pi * r^2

OpenStudy (xmissalycatx):

Close.. \[V=\Pi*r ^{2}*h\]

OpenStudy (theprofessor35):

Ohhh, so....now what do we do?

OpenStudy (xmissalycatx):

Okay, plug in your values that you were given into the formula. You will have 2 formulas.

OpenStudy (theprofessor35):

radius of 6 cm and a height of 24.5 cm. v = 3.14 * 6^2 * 24.5 v = 3.14 * 36 * 24.5

OpenStudy (xmissalycatx):

I have to go D:

OpenStudy (theprofessor35):

Oh okay. Bye! Thanks for starting me off! :DD